# Cost-Driven Representation Learning for Linear Quadratic Gaussian Control: Part I

### v3

@@ 4.6 Proof of Theorem 1
where $\overline{b}_{t}=\mathcal{O}(k)$, and $\overline{e}_{t}$ is a zero-mean subexponential random variable with $\|\overline{e}_{t}\|_{\psi_{1}}=\mathcal{O}(kd_{x}^{1/2})$. Then, it is clear that Algorithm 2 recovers latent states $z^{\ast\prime}_{t}=M^{\ast\prime}_{t}h_{t}$, where $0\leq t\leq T$, by a combination of quadratic regression and low-rank approximate factorization. Below we drop the superscript prime for notational simplicity, but keep in mind that the optimal state representation function $(M^{\ast}_{t})_{t=0}^{T}$, the corresponding latent states $(z^{\ast}_{t})_{t=0}^{T}$, and the true latent system parameters $((A^{\ast}_{t},B^{\ast}_{t})_{t=0}^{T-1},(Q^{\ast}_{t})_{t=0}^{T})$ are all with respect to the normalized parameterization.
@@
By Wainwright (2019, Theorem 6.1), with probability at least 1p1-p, as long as $n\geq(t\vee 1)(d_{y}+d_{u})+\log(1/p)$,
@@
where $(i)$ is due to Marshall et al. (1979, H.1.i, Chapter 9) and that the minimum positive singular values of $\sum\nolimits_{i=1}^{n}\hat{M}_{t}h_{t}^{(i)}(h_{t}^{(i)})^{\top}\hat{M}_{t}^{\top}$ and $\hat{M}_{t}$ are both their $k$th singular value, where $k$ is the rank of $\hat{M}_{t}$, by considering the singular value decomposition of $\hat{M}_{t}$. By standard concentration in analyzing linear regression (Wainwright, 2019), with probability at least 1p1-p, as long as $n\geq a\log(1/p)$ for some absolute constant $a>0$, $\sigma_{\min}\big{(}\sum\nolimits_{i=1}^{n}h_{t}^{(i)}(h_{t}^{(i)})^{\top}\big{)}=\Omega(n)$. Hence,
@@
is a constant, and it suffices to consider the latent state space for studying the policy suboptimality gap. In the latent state space, action $u_{t}=\hat{K}_{t}\hat{M}_{t}h_{t}=\hat{K}_{t}(S_{t}z^{\ast}_{t}+\delta_{t})$, where $\delta_{t}:=(\hat{M}_{t}-S_{t}M^{\ast}_{t})h_{t}$ is a Gaussian noise vector correlated with $z^{\ast}_{t}$. Since $h_{0}=y_{0}=C^{\ast}_{0}z^{\ast}_{0}+C^{\ast}_{0}(x_{0}-z^{\ast}_{0})+v_{0}$, we have $\|\mathrm{Cov}(y_{0})\|_{2}=\mathcal{O}(\|\mathrm{Cov}(z^{\ast}_{0})\|_{2})$ and $\|\mathrm{Cov}(\delta_{0})\|_{2}=\mathcal{O}(\|\hat{M}_{0}-S_{0}M^{\ast}_{0}\|_{2}^{2}\|\mathrm{Cov}(z^{\ast}_{0})\|_{2})$. Define $\iota_{0}:=\|\hat{M}_{0}-S_{0}M^{\ast}_{0}\|_{2}$ and $\iota_{t}:=t^{1/2}(1+\max_{\tau\leq t-1}\|\hat{K}_{\tau}\|_{2})\|\hat{M}_{t}-S_{t}M^{\ast}_{t}\|_{2}$ for $1\leq t\leq T-1$. Below we use induction to show that as long as $(\|\hat{M}_{t}-S_{t}M^{\ast}_{t}\|_{2})_{0\leq t\leq T-1}$ are small enough, $\|\mathrm{Cov}(\delta_{t})\|_{2}\leq\iota_{t}^{2}\max_{\tau\leq t}\|\mathrm{Cov}(z^{\ast}_{\tau})\|_{2}$ for all $0\leq t\leq T-1$.
@@ Acknowledgment
YT, SS acknowledge partial support from NSF CCF-2112665 (TILOS AI Research Institute). KZ acknowledges the support from the Simons-Berkeley Research Fellowship, the U.S. Army Research Office grant W911NF-24-1-0085, and the NSF CAREER Award 2443704. SS and KZ mainly worked on this project while at MIT. The authors also thank Xiang Fu, Horia Mania, and Alexandre Megretski for helpful discussions.
@@
The distribution of $\|g\|_{2}$ is known as $\chi$ distribution, and we know that $\|\|g\|\|_{\psi_{2}}=c^{\prime}d^{1/4}$ for an absolute constant $c^{\prime}>0$ (see, e.g., Wainwright (2019)). Hence, $\|\|x\|\|_{\psi_{2}}\leq c^{\prime}d^{1/4}\|\Sigma_{x}\|_{2}^{1/2}$. Similarly, $\|\|y\|\|_{\psi_{2}}\leq c^{\prime}d^{1/4}\|\Sigma_{y}\|_{2}^{1/2}$. Since $\|\|x\|\|y\|\|_{\psi_{1}}\leq\|\|x\|\|_{\psi_{2}}\|\|y\|\|_{\psi_{2}}$ (see, e.g., Vershynin (2018, Lemma 2.7.7)), we have
@@
$\leq\;$ $\|D_{x}^{\top}D_{x}\|_{2}+\|D_{y}^{\top}D_{y}\|_{2}=\|\mathbb{E}[xx^{\top}]\|_{2}+\|\mathbb{E}[yy^{\top}]\|_{2},$
completing the proof. ∎

### v4

@@ 4.6 Proof of Theorem 1
where $\overline{b}_{t}=\mathcal{O}(k)$, and $\overline{e}_{t}$ is a zero-mean subexponential random variable with $\|\overline{e}_{t}\|_{\psi_{1}}=\mathcal{O}(kd_{x}^{1/2})$.
Then, it is clear that Algorithm 2 recovers latent states $z^{\ast\prime}_{t}=M^{\ast\prime}_{t}h_{t}$, where $0\leq t\leq T$, by a combination of quadratic regression and low-rank approximate factorization. Below we drop the superscript prime for notational simplicity, but keep in mind that the optimal state representation function $(M^{\ast}_{t})_{t=0}^{T}$, the corresponding latent states $(z^{\ast}_{t})_{t=0}^{T}$, and the true latent system parameters $((A^{\ast}_{t},B^{\ast}_{t})_{t=0}^{T-1},(Q^{\ast}_{t})_{t=0}^{T})$ are all with respect to the normalized parameterization.
@@
By (Wainwright, 2019, Theorem 6.1), with probability at least 1p1-p, as long as $n\geq(t\vee 1)(d_{y}+d_{u})+\log(1/p)$,
@@
where $(i)$ is due to (Marshall et al., 1979, H.1.i, Chapter 9) and that the minimum positive singular values of $\sum\nolimits_{i=1}^{n}\hat{M}_{t}h_{t}^{(i)}(h_{t}^{(i)})^{\top}\hat{M}_{t}^{\top}$ and $\hat{M}_{t}$ are both their $k$th singular value, where $k$ is the rank of $\hat{M}_{t}$, by considering the singular value decomposition of $\hat{M}_{t}$. By standard concentration in analyzing linear regression (Wainwright, 2019), with probability at least 1p1-p, as long as $n\geq a\log(1/p)$ for some absolute constant $a>0$, $\sigma_{\min}\big(\sum\nolimits_{i=1}^{n}h_{t}^{(i)}(h_{t}^{(i)})^{\top}\big)=\Omega(n)$. Hence,
@@
is a constant, and it suffices to consider the latent state space for studying the policy suboptimality gap.
In the latent state space, action $u_{t}=\hat{K}_{t}\hat{M}_{t}h_{t}=\hat{K}_{t}(S_{t}z^{\ast}_{t}+\delta_{t})$, where $\delta_{t}:=(\hat{M}_{t}-S_{t}M^{\ast}_{t})h_{t}$ is a Gaussian noise vector correlated with $z^{\ast}_{t}$. Since $h_{0}=y_{0}=C^{\ast}_{0}z^{\ast}_{0}+C^{\ast}_{0}(x_{0}-z^{\ast}_{0})+v_{0}$, we have $\|\mathrm{Cov}(y_{0})\|_{2}=\mathcal{O}(\|\mathrm{Cov}(z^{\ast}_{0})\|_{2})$ and $\|\mathrm{Cov}(\delta_{0})\|_{2}=\mathcal{O}(\|\hat{M}_{0}-S_{0}M^{\ast}_{0}\|_{2}^{2}\|\mathrm{Cov}(z^{\ast}_{0})\|_{2})$. Define $\iota_{0}:=\|\hat{M}_{0}-S_{0}M^{\ast}_{0}\|_{2}$ and $\iota_{t}:=t^{1/2}(1+\max_{\tau\leq t-1}\|\hat{K}_{\tau}\|_{2})\|\hat{M}_{t}-S_{t}M^{\ast}_{t}\|_{2}$ for $1\leq t\leq T-1$. Below we use induction to show that as long as $(\|\hat{M}_{t}-S_{t}M^{\ast}_{t}\|_{2})_{0\leq t\leq T-1}$ are small enough, $\|\mathrm{Cov}(\delta_{t})\|_{2}\leq\iota_{t}^{2}\max_{\tau\leq t}\|\mathrm{Cov}(z^{\ast}_{\tau})\|_{2}$ for all $0\leq t\leq T-1$.
@@ Acknowledgment
YT, SS acknowledge partial support from NSF CCF-2112665 (TILOS AI Research Institute). KZ acknowledges the support from the Simons-Berkeley Research Fellowship, the U.S. Army Research Office grant W911NF-24-1-0085, and the NSF CAREER Award 2443704. The authors also thank Xiang Fu, Horia Mania, and Alexandre Megretski for helpful discussions.
@@
The distribution of $\|g\|_{2}$ is known as $\chi$ distribution, and we know that $\|\|g\|\|_{\psi_{2}}=c^{\prime}d^{1/4}$ for an absolute constant $c^{\prime}>0$ (see, e.g., (Wainwright, 2019)). Hence, $\|\|x\|\|_{\psi_{2}}\leq c^{\prime}d^{1/4}\|\Sigma_{x}\|_{2}^{1/2}$.
Similarly, $\|\|y\|\|_{\psi_{2}}\leq c^{\prime}d^{1/4}\|\Sigma_{y}\|_{2}^{1/2}$. Since $\|\|x\|\|y\|\|_{\psi_{1}}\leq\|\|x\|\|_{\psi_{2}}\|\|y\|\|_{\psi_{2}}$ (see, e.g., (Vershynin, 2018, Lemma 2.7.7)), we have
@@
$\leq\;$ $\|D_{x}^{\top}D_{x}\|_{2}+\|D_{y}^{\top}D_{y}\|_{2}=\|\mathbb{E}[xx^{\top}]\|_{2}+\|\mathbb{E}[yy^{\top}]\|_{2}.$